OpenStudy (anonymous):

Multiply. (5x+8y)(5x-8y)

OpenStudy (anonymous):

Distribute the numbers by multiplying them

OpenStudy (anonymous):

\[\left( ax+by \right)\left( ax-by \right)=a^2x^2-b^2y^2\]

OpenStudy (anonymous):

yessir